OpenStudy (anonymous):

use formula for compound interest to find A. A=p(1+r)n,p=$40000,r=0.15, n=4 years. end of 4 years A is what

OpenStudy (anonymous):

A. A=p(1+r)n, plug values of p n r and findA

OpenStudy (anonymous):

a=40000(1+.15)4 a=40000(1.15)4 multiply

OpenStudy (anonymous):

what do I mulp

OpenStudy (anonymous):

forget the four on the end. mulp 40000 and 1.15

OpenStudy (anonymous):

46000

OpenStudy (anonymous):

what he said

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is thatcorect

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks